OpenStudy (anonymous):

Solve the system by either substituition or elimination: 2x+5y=-22 5x-7y=23

OpenStudy (anonymous):

solve it

OpenStudy (anonymous):

Thanks :)

OpenStudy (saifoo.khan):

PLease hit good answer for imran. thank you

OpenStudy (anonymous):

To solve, you need to eliminate one of the variables. To eliminate one of the variables, you need to multiply the equations with a number that allows you to eliminate one of the variables.

OpenStudy (anonymous):

Thank you GT...much more helpful

OpenStudy (anonymous):

For example, multiplying the first equation by 7 and second equation by 5 will allow you to get rid of y.

OpenStudy (anonymous):

So 14x + 35y = -154 25x - 35y = -115

OpenStudy (karatechopper):

ok so u claick good answer then say hi and then say goodbye and then when u click good answer for me b sure to smile. o and to solve the problem u do this then that. and then if u still cant solve it then push the problem aside and yell at it saying "Y R U SO HARD TO FIGURE OUT!!!" then walk away...;)

OpenStudy (anonymous):

Right. Now, add the two equations. you will get an equation in just x. Solve for x. Then, substitute that value of x into one of the equations to find the value of y.

OpenStudy (anonymous):

It is +115 not -115.

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Subtract the two lines?

OpenStudy (anonymous):

I call it add. When you add 35y and -35y, you eliminate y. You can think of subtracting if that helps.

OpenStudy (anonymous):

x=-1?

OpenStudy (mathteacher1729):

2x+5y=-22 multiply this entire equation by 5 5x-7y=23 multiply this entire equation by 2 10x + 25y = -110 10x - 14y = 46 Subtract this entire equation from the first equation. (or, equivalently, multiply it by negative one) 10x + 25y = -110 -10x + 14y = -46 ============== 0x + 39y = -156 divide both sides by 39 to find y. Once you have y, substitute it back into either of the original equations to solve for x. :)

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

You got it. Good job.

OpenStudy (anonymous):

And then like MT said, plug it in and solve for y right?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

y=-4